OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (10, –5) and a radius of 11?

OpenStudy (anonymous):

\[(x-10)^2+(y+5)^2=121\]

OpenStudy (anonymous):

General formula: center (a,b) with radius r \[(x-a)^2+(y-b)^2=r^2\]